40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  Cryo Lab eLog  Not logged in ELOG logo
Entry  Thu Apr 28 11:19:24 2022, shruti, Update, PSOMA, Switch current driver to custom low noise 
    Reply  Thu Apr 28 14:44:28 2022, aaron, Update, PSOMA, Switch current driver to custom low noise fitted_full.pdfCalibratedNoiseBudget.pdf
       Reply  Fri Apr 29 11:21:51 2022, shruti, Update, PSOMA, Switch current driver to custom low noise CalibratedNoiseBudget.pdflaser_noise_comparison.pdf
Message ID: 2885     Entry time: Thu Apr 28 14:44:28 2022     In reply to: 2884     Reply to this: 2886
Author: aaron 
Type: Update 
Category: PSOMA 
Subject: Switch current driver to custom low noise 

[shruti, aaron]

PDH calibration

Later in the afternoon, we again measured the PDH error signal sweep driving with 2V from the function generator at 1.3 kHz. I saved the trace in cryo_lab/data/PDH/cal/220428_PDHcal.csv.

The best fit parameters are

  • Slope of the error signal is 0.33 V/MHz, or 2.99 MHz/V
  • Fitted cavity length is 30.1 cm, and the FSR is 998 MHz
  • Fitted transmission of the cavity is 1.0%
  • The relative RF and LO phase at the mixer is 198 degrees (nearly out of phase, off by 18 degrees)

Dark noise measurement

We blocked the beam and measured the noise on the PDH error signal using an SR785. We sent the PDH error signal from the error mon point on the LB servo box (same location we used for the calibration above) directly to the spectrum analyzer.

The new curve for photodiode dark noise is at 1 Hz/rtHz above a few kHz, and the curve lies below our previous estimate for the frequency noise of the laser -- progress!

Calibration pipeline

I'm making a file calpar.yml in our scripts/calibrate_noise directory. This file will hold the key variables for generating the noise budget, including file names and paths, cavity parameters, and which PDH calibration to use for various measurements.

Beat note

Shruti measured the frequency noise on the beat note in the same manner as above. More details and results to follow.

Shot noise

We didn't measure the shot noise, but we estimated to see the shot noise above the observed photodiode dark noise would require about 3 mW of power at 1550nm on the photodiode. We can then scale the shot noise by the actual power level with the cavity locked.

Attachment 1: fitted_full.pdf  59 kB  Uploaded Thu Apr 28 16:51:21 2022  | Hide | Hide all
fitted_full.pdf
Attachment 2: CalibratedNoiseBudget.pdf  204 kB  Uploaded Thu Apr 28 18:22:14 2022  | Hide | Hide all
CalibratedNoiseBudget.pdf
ELOG V3.1.3-